OpenStudy (anonymous):

What is the measure of each exterior angle in a regular 10-sided polygon? A. 18° B. 180° C. 360° D. 36°

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Thankss (:

OpenStudy (anonymous):

How is it 36?

OpenStudy (anonymous):

like for example a square = 4 sides.. 360/4 = 90 so 360/10 = 36

OpenStudy (anonymous):

bu the interior angles add up to more than just 360.

OpenStudy (anonymous):

what are the other angles? if its a decagon, which is a 10-gon.. then all the angles are the same..

OpenStudy (anonymous):

a square has 4 sides.. make a circle with and each quarter of the circle you put a point... you will have 4 points, you connect the points and you have a square.. 360/4= 90

OpenStudy (anonymous):

ok